OpenStudy (anonymous):

if sinθ=3/5 and θ terminates in the first quadrant, find the exact value of cos2θ

OpenStudy (anonymous):

i'm gonna use x instead of theta... since x is in the first quadrant, cosx=4/5... (use this is a 3-4-5 triangle) so using the double angle formula: cos2x = cos^2(x) - sin^2(x) = (4/5)^2 - (3/5)^2 can you take it from here?

OpenStudy (anonymous):

@viictor Do you still need help?

OpenStudy (anonymous):

7/25?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Do you up why -> cosx=4/5 ?

OpenStudy (anonymous):

Do you know why?

OpenStudy (anonymous):

its a 3 4 5

OpenStudy (anonymous):

cos2x = cos^2(x) - sin^2(x) = (4/5)^2 - (3/5)^2 So you okie here, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So you're fine with your result :)

OpenStudy (anonymous):

is it 7/25?

OpenStudy (anonymous):

(16 - 9 )/ 25 = 7/ 25

OpenStudy (anonymous):

so yes thanks

OpenStudy (anonymous):

The key is knowing the trig, formula!

OpenStudy (anonymous):

theyres to many

OpenStudy (anonymous):

Yup. there's only one way to make yourself better is practice more, like I'm doing now :)

OpenStudy (anonymous):

what about this one If sinθ=-1/4 and θ terminates in third quad., find the exact value of sin 2θ

OpenStudy (anonymous):

Could you open the new post, pls!

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

Start a new post , new page!